OpenStudy (anonymous):

Solve the equation. Use complete sentences to explain how you solved this equation.

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 } - \frac{ 1 }{ x-4 } = \frac{ -2 }{ 2x-8 }\]